OpenStudy (anonymous):

Match the degree with each polynomial function or if it is not a polynomial function x4 + 3x2 -5x 5x3y2 + 13x4y3-17x3y4 5x10 + 13x4 -17x3 x7 + 3 5x2 - 3x5 -5x3 this are the answers 10 4 5 9 7 not a polynomial function in 1variable please help i dont undersstand

OpenStudy (anonymous):

???

OpenStudy (shane_b):

The degree of a polynomial is determined by the highest degree of the terms in the expression. Look at the variables in each term and add up the exponents to get the degree of that term. Some examples: 3x^2 has a degree of 2 since x is raised to 2. 5x^2y^3 has a degree of 5 (2+3 = 5) 5x has a degree of 1 (since x=x^1) I'll also do one your problems for you: 5x3y2 + 13x4y3-17x3y4 The second and third terms have the highest degree in the expression...both have a degree of 7 (3 + 4= 7). The first term in this problem is 5...which is less than 7...so the entire expression has a degree of 7 since that's the highest degree of the individual terms. I hope that's clear :)

OpenStudy (anonymous):

thanks...